# Neuromorphic Quantum Neural Networks with Tunnel-Diode Activation Functions

### v1

@@ Abstract
The mathematical complexity and high dimensionality of neural networks hinder the training and deployment of machine learning (ML) systems while also requiring substantial computational resources. This fundamental limitation drives ML research, particularly in the exploration of alternative neural network architectures that integrate novel building blocks, such as advanced activation functions. Tunnel diodes are well-known electronic components that utilise the physical effect of quantum tunnelling (QT). Here, we propose using the current–voltage characteristic of a tunnel diode as a novel, physics-based activation function for neural networks. We demonstrate that the tunnel-diode activation function (TDAF) outperforms traditional activation functions in terms of accuracy and loss during both training and evaluation. We also highlight its potential for implementation in electronic circuits suited to developing neuromorphic, quantum-inspired AI systems capable of operating in environments not suitable for qubit-based quantum computing hardware.
@@ I Introduction
The fields of machine learning (ML) and neuromorphic computing (NC) are rapidly evolving, with increasing demands for energy-efficient and biologically plausible computational models [1, 2]. Algorithmic neural network models rely on activation functions [3]—mathematical functions applied to the output of a neuron—that enable deep learning architectures to learn complex patterns in the input data [4]. The nonlinearity of activation functions is crucial for modelling intricate relationships between inputs and outputs, enabling the neural network to generalise beyond the training data [5, 3, 4]. Without nonlinearity, neural networks reduce to linear models, limiting their ability to capture and learn hierarchical representations found in real-world data [3]. However, commonly used activation functions such as ReLU, sigmoid and tanh exhibit limitations, particularly in deep learning, due to issues such as vanishing gradients and energy inefficiency [4]. These challenges hinder learning in deep networks, necessitating alternative solutions that offer both efficient computation and reduced power consumption [4, 6, 7].
Physics, particularly quantum mechanics, has profoundly influenced ML, inspiring the development of novel algorithms that exploit physical principles to enhance computational efficiency and reduce power consumption [1, 2, 8, 9]. A notable example is quantum reservoir computing (RC) [10, 11, 12], a subclass of neuromorphic and unconventional computing systems [13] that employs the intrinsic dynamics of physical systems to enable energy-efficient ML solutions, particularly for predicting highly nonlinear and chaotic time series, as well as for pattern recognition tasks [11, 12]. Concurrently, physics-informed neural networks integrate fundamental physical laws—expressed through partial differential equations—directly into the learning process [14]. This approach ensures that models remain consistent with established physical principles while requiring fewer data points compared to traditional deep learning methods.
Quantum neural networks (QNNs) are computational models that integrate quantum mechanics with traditional neural network architectures used in ML [15, 16]. Modern research on QNNs focuses on integrating classical artificial neural networks—widely used in ML and computer vision [4]—with the unique advantages of quantum information processing [17, 18].
QNNs and NC systems share the goal of replicating cognitive processes observed in biological brains [19, 20, 21, 8, 9, 22]. Both QNNs and NC systems aim to overcome the limitations of conventional computing, which relies on deterministic logic and sequential processing. NC systems address these challenges through event-driven architectures and parallel processing, mimicking the function of biological neurons and synapses [19, 20, 21].
@@
Quantum tunnelling (QT) has been utilised in tunnel [23] and resonant-tunnelling diodes [24], with certain NC architectures exploiting negative differential resistance [25, 22], a hallmark of tunnelling diodes. QT-based devices, including tunnel diodes, exhibit significantly lower power consumption than traditional integrated circuits [26], also exhibiting a number of characteristics essential for applications in challenging environments such as space [27]. Additionally, QT has facilitated the development of scanning tunnelling microscopy (STM) [28] and has also been experimentally observed in quantum dots [25], which serve as essential components of quantum NC systems [10].
Recent advancements in quantum computing have increasingly employed QT in quantum annealing and NC systems [29]. In quantum annealing, tunnelling enables systems to traverse potential energy barriers, aiding the discovery of optimal solutions in combinatorial optimisation problems [29]. In other classes of NC systems, QT mechanisms have been integrated to emulate the architecture of the brain [30]. Finally, QT also plays a critical role in quantum biology, influencing enzymatic reactions and energy transfer in photosynthesis [31].
The mathematical expressions governing the probability of QT through a potential barrier exhibit rich nonlinear properties [32] and semiconductor diodes serve as a key element of NC systems [33]. Therefore, in this paper we propose and theoretically demonstrate that the current–voltage (I–V) characteristic of a tunnel diode can serve as a novel nonlinear activation function in neural networks (Fig. 1). Unlike conventional activation functions, which are often heuristic and lack direct physical interpretations, the intrinsic nonlinearity of the tunnel diode arises from well-defined quantum mechanical principles [34, 35]. By systematically comparing traditional activation functions used in ML, we show that the neural network implementing the tunnel-diode activation function (TDAF) not only exhibits superior performance in terms of convergence, accuracy and expressiveness of the model but also introduces an energy-efficient approach to neuromorphic computing by extending prior research work [32]. This quantum-inspired activation function opens new possibilities for designing more biologically plausible and computationally efficient neural architectures, further bridging the gap between semiconductor physics and quantum AI systems.
@@ II-A Traditional Activation Functions
For example, sigmoid $f(x)=\dfrac{1}{1+e^{-x}}$ and hyperbolic tangent (tanh) $f(x)=\dfrac{e^{x}-e^{-x}}{e^{x}+e^{-x}}$ functions [19, 36, 4] compress inputs into a fixed range $(0,1)$ and $(-1,1)$, respectively. However, while useful for probability estimation and feature scaling, these functions suffer from saturation, where large or small inputs lead to near-zero gradients, hindering weight updates in multilayer networks [4].
Therefore, one of the most widely used activation functions is the Rectified Linear Unit (ReLU), defined as $f(x)=\max(0,x)$ [4]. ReLU is computationally efficient and helps mitigate some of the vanishing gradient problems associated with sigmoid and tanh functions. However, ReLU is not fully immune to gradient-related issues, since neurons based on this function can become inactive when they consistently receive negative inputs, which prevents further learning [4]. To address this limitation, variants such as Leaky ReLU and Parametric ReLU have been proposed, which introduce small negative slopes for negative inputs to maintain gradient flow and prevent inactive neurons [6]. Other functions such as Swish $f(x)=x\cdot\sigma(x)$, where $\sigma(x)$ is the sigmoid function, and GELU have been introduced to balance smoothness and gradient propagation [7]. Commonly used in transformer models, these activation functions offer improved expressiveness but remain not universally optimal [7].
@@ II-B Tunnel-Diode Activation Function
A tunnel diode exploits the effects of QT, enabling charge carriers to pass through an energy barrier rather than overcoming it thermally [23]. Unlike conventional p-n junctions, tunnel diodes are heavily doped, reducing the depletion width to just a few nanometres [35]. This extreme doping causes the conduction band of the n-region and the valence band of the p-region to become energetically aligned, enabling electrons to tunnel directly through the barrier.
At zero bias and small forward bias, electrons tunnel from the conduction band of the n-region to the valence band of the p-region, resulting in a sharp increase in current. As voltage increases further, band misalignment reduces the probability of QT, causing a drop in current—a phenomenon known as negative differential resistance. At even higher voltages, QT ceases entirely and the diode behaves like a standard p-n junction. This transport mechanism enables tunnel diodes to operate at high speeds and low power levels, making them valuable for high-frequency and microwave applications.
Mathematically, the current-voltage (I–V) characteristic curve of a tunnel diode is a highly nonlinear function, primarily due to a negative differential resistance region, and it is given by
@@
The parameters are defined as $\alpha=\dfrac{q(b-c)}{k_{B}T}$, $\eta=\dfrac{qn_{1}}{k_{B}T}$ and $\gamma=\dfrac{qn_{2}}{k_{B}T}$, where $q$ is the electron charge, $k_{B}$ is Boltzmann’s constant, $T$ is the temperature, $V$ is the voltage, and $a$, $b$, $c$, $d$, $n_{1}$ and $n_{2}$ are parameters characterising the physical system. In this work, we use $T=300$ K, $a=0.0039$ A, $b=0.5$ V, $c=0.0874$ V, $d=0.0073$ V, $n_{1}=0.0352$, $n_{2}=0.0031$ and $h=0.0367$ A, with all as per Ref. [37] except $b$, which has been scaled by a factor of $10$ to explore the nonlinear behaviour of TDAF.
@@
When compared to ReLU and sigmoid (Fig. 2a, c), TDAF, from an ML perspective, exhibits two main regions: a nominally weakly nonlinear region for negative inputs which corresponds to the first positive differential resistance region of the I-V characteristic, and a nominally highly nonlinear region for positive inputs which corresponds to the negative differential resistance region and the second positive differential resistance region of the I-V characteristic. Upon comparing the derivative plots of ReLU and TDAF (Fig. 2b, f), it is reasonable to expect that these two functions will exhibit similar ML behaviour. Additionally, we argue that TDAF exhibits a greater degree of nonlinearity compared to ReLU and sigmoid activation functions, as demonstrated below.
II-C Analysis of Nonlinear Properties of TDAF
Figure 3: Fourier spectra of the outputs of (b) ReLU, (c) sigmoid and (d) TDAF functions activated by a sinusoidal wave signal at a frequency of 16 Hz. Panel (a) shows the spectrum of the reference output produced by the linear identity function.
Analysing the degree of nonlinearity is a non-trivial task, as no standard metrics exist, and nonlinearity itself arises from diverse physical and mathematical processes. For example, one method to compare the nonlinearity of activation functions is to analyse their Taylor series expansions and compare the magnitudes of the higher-order terms. However, the Taylor series is an approximation and calculating its coefficients can be non-trivial, especially for artificial functions such as ReLU.
Consequently, a previous study [39] suggested that two nonlinear physical phenomena can be compared by examining the Fourier spectra of a linear signal processed by the nonlinear systems of interest. Following this approach, we analyse the responses of the ReLU, sigmoid and TDAF activation functions to a purely sinusoidal wave signal at a frequency of 16 Hz. For the sake of comparison, we also apply the sinusoidal signal to a linear identity function, producing a Fourier spectrum with a sole peak at 16 Hz.
As shown in Fig. 3, the response of ReLU, sigmoid and TDAF to the sinusoidal waves results in the nonlinear generation of higher-order harmonics. According to [39], the strength of the nonlinearity can be quantified using the magnitude of the harmonic peaks and the total number of harmonics produced by the nonlinear process. Focusing on the latter criterion, we can see that ReLU produces the peak at the second-harmonic frequency 32 Hz and then at the fourth, 64 Hz, and the sixth, 96 Hz, harmonics. In turn, sigmoid produces the third, 48 Hz, and fifth, 80 Hz, peaks only. On the contrary, TDAF generates strong peaks at both odd and even harmonic frequencies. Within the theoretical framework used in this analysis [39], this result suggests that the TDAF exhibits the strongest nonlinearity among the three analysed functions. We also observe that both TDAF and ReLU exhibit DC components in their respective spectra, which further supports our earlier assumption of their similar responses.
The selection of a simple sinusoidal input signal for analysing the spectra of the activation functions carries significant meaning from an ML and NC perspective. It was demonstrated that a traditional RC algorithm, which can, in principle, employ any standard neural activation function [19], is mathematically equivalent to a model that uses both linear weights of an input dataset and its nonlinear (quadratic, cubic and so on) functionals [40]. It follows from Ref. [40] that the computational efficiency of the model can be increased by constructing an algorithm or using a suitable, either theoretical or real-life physical, dynamical system that is highly nonlinear enough to expand the input data over many higher-order nonlinear functionals. This assertion was confirmed experimentally using a fluid-mechanical dynamical system employed as a neuromorphic computer [41]. In light of this discussion, TDAF outperforms both ReLU and sigmoid, as its Fourier spectrum contains all nonlinear harmonics, with their respective magnitudes being stronger, thus enabling better generation of nonlinear functionals of the input data and resulting in higher computational efficiency.
II-D Neural Networks
Feedforward neural networks (FNNs)—the foundational model considered in this paper—typically consist of an input layer, $L$ hidden layers and an output layer [42]. The $l$th hidden layer, containing $N_{l}$ neurons, takes an input vector, $x^{(l-1)}$, from the previous layer and applies a linear transformation
@@ II-D Neural Networks
where $W^{(l)}\in M^{N_{l-1}\times N_{l}}$ and $b^{(l)}\in\mathbb{R}^{N_{l}}$ are the weight matrix and bias vector of layer $l$ that are applied to input data received from layer $l-1$. An activation function, $\sigma^{(l)}$ is then applied to the output to obtain the input to the next layer $x^{(l)}=\sigma^{(l)}\left(a^{(l)}\right)$, where $a^{(l)}=F\left(x^{(l-1)}\right)$ [43, 44]. The neural network $\mathcal{N}$ is then the composition of the functions comprising these layers (see Fig. 1 for an illustration)
@@
Thus, nonlinearity is introduced into the neural network through the activation functions $\sigma_{1},\dots,\sigma_{L+1}$, which, in the framework of an individual model, are typically all the same function, $\sigma_{i}=\sigma:\mathbb{R}\to\mathcal{R},\quad\forall i$ [5, 3, 4]. FNNs employing nonlinear activation functions have been shown to be universal function approximators [45]. For many common activation functions, the range, $\mathcal{R}$, is $[0,1]$ or $[-1,1]$.
@@
II-E Numerical Experiments and Datasets
Figure 1 illustrates the general feed-forward neural network architecture used in this paper, with the inset specifically showing how TDAF is incorporated into the model. Each network consists of three layers: an input layer, a single hidden layer and an output layer. Three topologically identical neural networks were implemented in PyTorch, with the only distinction being the activation function applied after the hidden layer: ReLU, sigmoid or TDAF.
@@ II-E Numerical Experiments and Datasets
Additionally, to prevent extreme numerical values from entering Eq. (1), the input was clamped between $-10$ and $10$ (the choice of clamping limits depends on the specific data and task). The output was normalised between 00 and $1$. The corresponding source code is available at [49].

### v2

# Full-Precision and Ternarised Neural Networks with Tunnel-Diode Activation Functions: Computing and Physics Perspectives

@@ Abstract
The mathematical complexity and high dimensionality of neural networks slow both training and deployment, demanding heavy computational resources. This has driven the search for alternative architectures built from novel components, including new activation functions. Taking a different approach from state-of-the-art neural and neuromorphic computational systems, we employ the current–voltage characteristic of a tunnel diode as a quantum physics-based activation function for deep networks. This tunnel-diode activation function (TDAF) outperforms standard activations in deep architectures, delivering lower loss and higher accuracy in both training and evaluation. We also highlight its promise for implementation in electronic hardware aimed at neuromorphic, ternarised and energy efficient AI systems. Speaking broadly, our work lays a solid foundation for a new bridge between machine learning, semiconductor electronics and quantum physics—bringing together quantum tunnelling, a phenomenon recognised in six Nobel Prizes (including the 2025 award), and contemporary AI research.
@@ I Introduction
The field of machine learning is evolving quickly, driven by growing demand for energy-efficient and biologically plausible models. Neural networks depend on activation functions [1]—nonlinear transformations applied to each neuron’s output—that enable deep architectures to learn complex structure in data [2, 1, 3]. This nonlinearity is essential: without it, a network collapses to a simple linear model and cannot capture the layered representations present in real-world signals [1]. Yet widely used activations such as ReLU, sigmoid and tanh face well-known limitations in deep learning, including vanishing gradients and poor energy efficiency [1, 4, 3]. These problems restrict scalability and slow learning in deep networks, underscoring the need for alternative activation mechanisms that deliver robust performance with lower computational and power costs [3, 5, 6].
Physics, particularly quantum mechanics, has profoundly influenced machine learning, inspiring the development of novel algorithms that exploit physical principles to enhance computational efficiency and reduce power consumption [7, 8, 9, 10]. A notable example is quantum reservoir computing [11, 12, 13], a subclass of neuromorphic and unconventional computing systems [14] that employs the intrinsic dynamics of physical systems to enable energy-efficient machine learning solutions, particularly for predicting highly nonlinear and chaotic time series, as well as for pattern recognition tasks [12, 13]. Concurrently, spiking resonant tunnelling-diode neurons and neural networks [15, 16, 17, 18, 19] are gaining attention for their ability to reproduce key biological neuronal behaviours, including excitable spiking and refractoriness, while offering high operating speeds, energy efficiency and compact device footprints.
@@
In this paper, we take a fundamentally different path from neuromorphic spiking systems [15, 16, 17, 18, 19]—and, indeed, from any computational architecture that employs tunnel diodes or related quantum electronic devices [20]. Instead, we use the current–voltage characteristic of a tunnel diode as a quantum-physics-based activation function for deep neural networks.
Although our work ultimately converges with ongoing efforts to develop more efficient neuromorphic neural network systems employing tunnel diodes [15, 16, 17, 18, 19], its conceptual foundations arise from a completely different direction. In our case, the idea of employing quantum-tunnelling (QT) electron devices [21, 22] within a neural network originates in research on quantum cognition theory [23, 24, 25, 26], where quantum statistics has been shown to substantially outperform classical models [27] in explaining human cognitive behaviour, perception and decision-making [28, 29, 30].
Beyond the intriguing fact that QT, the defining physical effect underpinning tunnel diodes, offers a novel pathway for modelling the complexities of human cognition, QT has been employed in neuromorphic architectures that exploit negative differential resistance [31, 32]. QT-based devices also provide substantially lower power consumption than traditional integrated circuits [33] and possess properties essential for operation in harsh environments such as space [34]. QT underlies the operation of scanning tunnelling microscopy [35] and has been observed in quantum dots [31], which themselves constitute key building blocks in neuromorphic systems [11].
Recent advances in quantum computing further incorporate QT in quantum annealing and neuromorphic architectures [36], where tunnelling enables systems to traverse energy barriers and identify optimal solutions in combinatorial optimisation tasks [36]. QT-based mechanisms have also been integrated into computational models inspired by brain-like architectures [37] and QT plays a recognised role in quantum biology, influencing enzymatic processes and energy transfer in photosynthesis [38].
The mathematical expressions governing the probability of QT through a potential barrier exhibit rich nonlinear properties [28] and semiconductor diodes serve as a key element of NC systems [39]. Therefore, in this paper we propose and theoretically demonstrate that the current–voltage (IVI–V) characteristic of a tunnel diode can serve as a novel nonlinear activation function in neural networks (Fig. 1). Unlike conventional activation functions, which are often heuristic and lack direct physical interpretations, the intrinsic nonlinearity of the tunnel diode arises from well-defined quantum mechanical principles [40, 41]. We show that the neural network implementing the tunnel-diode activation function (TDAF) not only exhibits superior performance in terms of convergence, accuracy and expressiveness of the model but also introduces an energy-efficient approach to neuromorphic computing by extending prior relevant research work on neural networks with ternarised weights and activation functions [42, 43].
Finally, our work establishes a solid foundation for a new bridge between machine learning, semiconductor electronics and quantum physics—uniting quantum tunnelling, a phenomenon recognised by six Nobel Prizes (including the 2025 award) [44], with contemporary AI research.
@@ II-A Traditional Activation Functions
For example, sigmoid $f(x)=\dfrac{1}{1+e^{-x}}$ and hyperbolic tangent (tanh) $f(x)=\dfrac{e^{x}-e^{-x}}{e^{x}+e^{-x}}$ functions [45, 46, 3] compress inputs into a fixed range $(0,1)$ and (1,1)(-1,1), respectively. However, while useful for probability estimation and feature scaling, these functions suffer from saturation, where large or small inputs lead to near-zero gradients, hindering weight updates in multilayer networks [3].
Therefore, one of the most widely used activation functions is the Rectified Linear Unit (ReLU), defined as $f(x)=\max(0,x)$ [3]. ReLU is computationally efficient and helps mitigate some of the vanishing gradient problems associated with sigmoid and tanh functions. However, ReLU is not fully immune to gradient-related issues, since neurons based on this function can become inactive when they consistently receive negative inputs, which prevents further learning [3].
To address this limitation, variants such as Leaky ReLU and Parametric ReLU have been proposed, which introduce small negative slopes for negative inputs to maintain gradient flow and prevent inactive neurons [5]. Other functions such as Swish $f(x)=x\cdot\sigma(x)$, where $\sigma(x)$ is the sigmoid function, and GELU have been introduced to balance smoothness and gradient propagation [6]. Commonly used in transformer models, these activation functions offer improved expressiveness but remain not universally optimal [6].
@@ II-B Tunnel-Diode Activation Function
A tunnel (Esaki) diode is a heavily doped pnp–n junction electron device whose IVI-V characteristic contains a region of negative differential resistance arising from quantum-mechanical tunnelling through a narrow potential barrier [21, 41]. Heavy doping reduces the depletion width to only a few nanometres, aligning the $n$-region conduction band with the $p$-region valence band and enabling electrons to tunnel directly through the junction. Under zero or small forward bias this band alignment produces a sharp increase in current. As the voltage increases, band misalignment lowers the tunnelling probability and the current falls, giving rise to the characteristic negative differential resistance region.
Resonant-tunnelling diodes form a related class of devices in which quantum wells and double potential barriers enable resonant transport. Although structurally different, Esaki diodes and resonant-tunnelling diodes share the same physical basis: voltage-dependent QT that yields a highly nonlinear transport response [41, 22, 47, 48]. While the present work employs a theoretical model originally developed for resonant-tunnelling diodes [15, 16], the discussion applies broadly to tunnel diodes. Crucially, a tunnel diode provides a strongly nonlinear, voltage-dependent effective resistance [10], making it well suited as a physics-based activation function.
Mathematically, the IVI-V characteristic curve of a tunnel diode is a highly nonlinear function, primarily due to a negative differential resistance region, and it is given by
@@
The parameters are defined as $\alpha=\frac{q(b-c)}{k_{B}T}$, $\eta=\frac{qn_{1}}{k_{B}T}$ and $\gamma=\frac{qn_{2}}{k_{B}T}$, where $q$ is the electron charge, $k_{B}$ is Boltzmann’s constant, $T$ is the temperature, $V$ is the voltage, and $a$, $b$, $c$, $d$, $n_{1}$ and $n_{2}$ are parameters characterising the physical system. In this work, we use $T=300$ K, $a=0.0039$ A, $b=0.5$ V, $c=0.0874$ V, $d=0.0073$ V, $n_{1}=0.0352$, $n_{2}=0.0031$ and $h=0.0367$ A, with all as per Ref. [15] except $b$, which has been scaled by a factor of $10$. These coefficients act as tunable parameters in the model and may be linked to relevant semiconductor physics quantities [49].
Thus, Eq. (1) provides an accurate fit to the IVI-V characteristics of standard tunnel diodes because its fitting parameters are free to take values not strictly tied to direct physical interpretation [49]. Other models of tunnel diode large-signal behaviour have also been proposed [50] and these can be substituted for Eq. (1) when analysing specialised or non-conventional diode structures.
@@
When compared to ReLU and sigmoid (Fig. 2a, c), TDAF, from a machine learning perspective, exhibits two main regions: a nominally weakly nonlinear region for negative inputs which corresponds to the first positive differential resistance region of the IVI-V characteristic, and a nominally highly nonlinear region for positive inputs which corresponds to the negative differential resistance region and the second positive differential resistance region of the I-V characteristic. Upon comparing the derivative plots of ReLU and TDAF (Fig. 2b, f), it is reasonable to expect that these two functions will exhibit similar machine learning behaviour. Additionally, we argue that TDAF exhibits a greater degree of nonlinearity compared to ReLU and sigmoid activation functions, as demonstrated below.
II-C Conventional Full-Precision Neural Network Model
Feedforward neural networks, the foundational computational model considered in this paper, typically consist of an input layer, $L$ hidden layers and an output layer [51]. The $l$th hidden layer, containing $N_{l}$ neurons, takes an input vector, $x^{(l-1)}$, from the previous layer and applies a linear transformation
@@ II-C Conventional Full-Precision Neural Network Model
where $W^{(l)}\in M^{N_{l-1}\times N_{l}}$ and $b^{(l)}\in\mathbb{R}^{N_{l}}$ are the weight matrix and bias vector of layer $l$ that are applied to input data received from layer l1l-1. An activation function, $\sigma^{(l)}$ is then applied to the output to obtain the input to the next layer $x^{(l)}=\sigma^{(l)}\left(a^{(l)}\right)$, where $a^{(l)}=F\left(x^{(l-1)}\right)$ [52, 53]. The neural network $\mathcal{N}$ is then the composition of the functions comprising these layers (see Fig. 1 for an illustration)
@@
Thus, nonlinearity is introduced into the neural network through the activation functions $\sigma_{1},\dots,\sigma_{L+1}$, which, in the framework of an individual model, are typically all the same function, $\sigma_{i}=\sigma:\mathbb{R}\to\mathcal{R},\quad\forall i$ [2, 1, 3]. Feedforward neural networks employing nonlinear activation functions have been shown to be universal function approximators [54]. For many common activation functions, the range, $\mathcal{R}$, is $[0,1]$ or [1,1][-1,1].
@@
II-D Numerical Experiments and Datasets
Figure 1 illustrates the general feed-forward neural network architecture used in this paper, with the inset specifically showing how TDAF is incorporated into the model. Each network consists of three layers: an input layer, a single hidden layer and an output layer. In the first set of numerical experiments, three topologically identical neural networks were implemented in PyTorch, with the only distinction being the activation function applied after the hidden layer: ReLU, sigmoid or TDAF.
@@ II-D Numerical Experiments and Datasets
Additionally, to prevent extreme numerical values from entering Eq. (1), the input was clamped between 10-10 and $10$ (the choice of clamping limits depends on the specific data and task). The output was normalised between 0 and $1$. The corresponding source code is available at [57].